(5 pts) An Excel worksheet table contains 50 equally spaced data values in the same column for the domain: 3. 3x < x < 4r. The left most domain value-3x is entered in the C3 cell . (2 pts) Calculate the increment used to equally space the domain values in the worksheet. (I pt) Then enter the exact increment value into the BI cell. (2 pts) If the domain values are contiguous in the same column (Column C), complete the expression used to compute the next domain data value in the C4 cell. The best possible answer is the answer that can be applied to the remaining column cells with the copy operation. . 1 Increment: Domain =c3 + 5 (more questions on the back)
Answer #1

Clearly, the increment value can be calculated as

\frac{4\pi - (-3\pi)}{50} = \frac{7\pi}{50}

You can enter this value in the B1 Cell.

Now, as to what shall you enter into the cell C4. Well, They have given half of the expression which includes the term C3. Since this should work for if you copy the command into a higher row value, you must construct the command keeping in mind which elements are supposed to be constant and which are supposed to be variable. Clearly, you can calculate the value of any row as the sum of the values of the previous row and the increment. Therefore, C4=C3+B1 . Similarly, C5=C4+B1. Since B1 is constant throughout while C3 becomes C4, therefore, you must write B1 as $B$1 in order to preserve its value during the copy operation. Thus, the command at the cell C4 would be =C3 + $B$1 .